# Incentives in Federated Learning with Heterogeneous Agents

### v1

@@ A.1 Effect of Data Composition on Agent-Specific Performance
For varying total data size $m$ and fraction $\lambda\in[0,1]$ from Agent 1, we construct datasets $D^{(m,\lambda)}$ comprising $\lambda m$ samples from Agent 1 and $(1-\lambda)m$ samples from Agent 2. Models trained on these datasets are evaluated separately on each agent’s test set. We use standard cross-entropy loss and the SOAP optimizer [34]. To capture performance variability, experiments are repeated 100 times with different random seeds. See Table˜1 for detailed parameters. We employed task-appropriate model architectures aligned with prior federated learning works [11, 10, 6]. For FEMNIST (image classification), the model is a lightweight CNN. For Shakespeare (next-character prediction), we use a recurrent neural network with an LSTM backbone.
@@
These experiments confirm the intuitive expectations about data volume and personalization, and quantify their effects. (i) For any fixed share $\lambda$, increasing the budget $m$ lowers both agents’ losses—more data helps everyone (Assumption˜2). (ii) For any fixed $m$, raising $\lambda$ cuts Agent 1’s loss: when the training set includes a higher proportion of one writer’s style, the model is better tuned to that style and thus generalizes better on that agent’s test set. These trends mirror prior personalization findings in federated learning and domain adaptation [9].
@@ Results.
Appendix B On Binary Payoffs
We model each agent’s payoff as $1$ if its PAC requirement holds and 0 otherwise. With this choice, each agent solves a cost–minimization problem subject to its threshold:
@@ Appendix B On Binary Payoffs
We believe this individual cost–minimization task reflects real-world scenarios, [^7] For example, hospitals may aim to minimize their data contribution while ensuring that the global model achieves at least $99\%$ accuracy on their local patient population, being indifferent between $99\%$ and $99.5\%$ once the target is met. and by standard Lagrange multiplier theory, there exists a multiplier $\lambda>0$ such that any optimal solution $m_{i}^{\star}$ to the constrained problem is also optimal for a penalized objective of the form
$\max_{m_{i}}\ a_{i}(\bm{m})-\lambda m_{i}$
Hence the binary case, which made the exposition lighter, is not a fundamental limitation but a modeling convenience that extends to a broader family of monotone utilities that trade off accuracy against cost with weight $\lambda$.
@@ D.1 Proof of the NP-hardness Result (Theorem˜2)
We reduce from the NP-complete Set Cover problem [35]: given a universe $U=\{u_{1},\dots,u_{n}\}$, a family of subsets $\mathcal{E}=\{E_{1},\dots,E_{r}\}$ with $\bigcup_{i\in[r]}E_{i}=U$, and an integer $m$, decide whether at most $m$ subsets cover $U$.
@@ D.2 Proof of the Approximation Guarantee (Theorem˜3)
In learning theory, for any distribution $\mathcal{D}$, a subset $\overline{\mathcal{H}}\subseteq\mathcal{H}$ is a $\gamma$-cover of $\mathcal{H}$ under $\mathcal{D}$ if, for every $h\in\mathcal{H}$, there exists $\overline{h}\in\overline{\mathcal{H}}$ such that $\mathcal{D}(\mathrm{DIS}(h,\overline{h}))\leq\gamma$. Haussler’s sphere-packing bound guarantees [36] the existence of such a cover with

### v2

@@ A.1 Effect of Data Composition on Agent-Specific Performance
For varying total data size $m$ and fraction $\lambda\in[0,1]$ from Agent 1, we construct datasets $D^{(m,\lambda)}$ comprising $\lambda m$ samples from Agent 1 and $(1-\lambda)m$ samples from Agent 2. Models trained on these datasets are evaluated separately on each agent’s test set. We use standard cross-entropy loss and the SOAP optimizer (Vyas et al., 2024). To capture performance variability, experiments are repeated 100 times with different random seeds. See Table˜1 for detailed parameters. We employed task-appropriate model architectures aligned with prior federated learning works (McMahan et al., 2017; Caldas et al., 2018; Murhekar et al., 2024). For FEMNIST (image classification), the model is a lightweight CNN. For Shakespeare (next-character prediction), we use a recurrent neural network with an LSTM backbone.
@@
These experiments confirm the intuitive expectations about data volume and personalization, and quantify their effects. (i) For any fixed share $\lambda$, increasing the budget $m$ lowers both agents’ losses—more data helps everyone (Assumption˜2). (ii) For any fixed $m$, raising $\lambda$ cuts Agent 1’s loss: when the training set includes a higher proportion of one writer’s style, the model is better tuned to that style and thus generalizes better on that agent’s test set. These trends mirror prior personalization findings in federated learning and domain adaptation (Hsu et al., 2019).
@@ Results.
Appendix B Binary Payoffs and Quasi-Linear Utilities
In this section we explain why our binary “meets target or not” payoff should be viewed as a convenient normalization rather than a restrictive assumption. We model each agent’s payoff as $1$ if its PAC requirement holds and 0 otherwise, but show that this threshold formulation is equivalent, via a standard Lagrangian/KKT argument, to optimizing a much broader class of quasi-linear utilities that trade off accuracy against cost. In particular, any agent with smooth, monotone accuracy benefits and convex costs behaves as if it had such a hard accuracy requirement, so the binary model compactly represents a wide family of underlying preferences.
Our agent‑level decision problem is a cost–minimization subject to meeting its accuracy requirement:
@@ Appendix B Binary Payoffs and Quasi-Linear Utilities
Under standard regularity conditions (feasibility of the constraint, monotonicity and differentiability of the underlying accuracy function, and convexity of the cost), the Karush–Kuhn–Tucker (KKT) conditions imply that any optimal solution to this constrained problem can be obtained by optimizing an equivalent unconstrained objective. Let $A_{i}(m_{i})$ denote a smooth, increasing performance metric for agent $i$ (for example, one minus error) such that $A_{i}(m_{i}^{\star})=1$ at the optimum, and let the per-sample cost be $c_{i}>0$. The Lagrangian for agent $i$’s problem is
$\mathcal{L}_{i}(m_{i},\lambda_{i})=c_{i}m_{i}+\lambda_{i}\big(1-A_{i}(m_{i})\big),$
with multiplier $\lambda_{i}\geq 0$ on the accuracy constraint. At any optimal solution $m_{i}^{\star}$, the KKT stationarity condition gives
$\frac{\partial\mathcal{L}_{i}}{\partial m_{i}}(m_{i}^{\star})=c_{i}-\lambda_{i}A_{i}^{\prime}(m_{i}^{\star})=0\quad\Rightarrow\quad A_{i}^{\prime}(m_{i}^{\star})=\frac{c_{i}}{\lambda_{i}}.$
In particular, there exists a multiplier $\lambda_{i}>0$ such that at the optimum $m_{i}^{\star}$, the marginal benefit of contributing data, $\mathcal{A}_{i}^{\prime}(m_{i}^{\star})$, equals the marginal cost $c_{i}$ scaled by $\lambda_{i}$. Intuitively, $\lambda_{i}$ serves to balance accuracy and cost at the threshold point. It is exactly the shadow price of the accuracy constraint: it measures how much the agent’s minimum cost would increase if the required accuracy threshold were raised slightly.
This yields a precise equivalence between the threshold and quasi-linear viewpoints. Consider the unconstrained quasi-linear utility
$U_{i}(m_{i})=A_{i}(m_{i})=\widetilde{\lambda}_{i}m_{i},$
where $\widetilde{\lambda}_{i}$ is the agent’s marginal disutility for each contributed sample (we can absorb $c_{i}$ into $\widetilde{\lambda}_{i}$ by scaling). The first-order condition for maximizing $U_{i}$ is $A_{i}^{\prime}(m_{i})=\widetilde{\lambda}_{i}$. Choosing $\widetilde{\lambda}_{i}=c_{i}/\lambda_{i}$ makes this condition identical to the KKT stationarity condition above, so $m_{i}^{\star}$ is also optimal for the unconstrained problem. In other words, the same $m_{i}^{\star}$ solves both “minimize cost subject to meeting the threshold” and “maximize smooth quasi-linear utility that trades off accuracy and cost at rate $\widetilde{\lambda}_{i}$.” The optimal stopping point for a satisficing agent under a continuous accuracy–benefit curve is therefore identical to the stopping point under the binary threshold model. Ex post, using a threshold utility does not exclude the optimal outcome that a more complex utility would produce: for any smooth, monotone accuracy–benefit function and convex cost, there exists a cost weight $\widetilde{\lambda}_{i}$ such that the agent behaves as if it had a hard requirement.
The above argument relies on $A_{i}(m_{i})$ being monotone (more data never hurts) and differentiable; if we further assume $A_{i}$ is concave (diminishing returns on accuracy), the correspondence between the threshold model and a smooth utility model becomes tight. Under concavity, the stopping point $m_{i}^{\star}$ is uniquely characterized by the balance of marginal benefit and cost, and this point is exactly where it is optimal to “just meet” the target and not contribute further. In that sense, any agent with a smooth, monotone benefit of accuracy and convex cost behaves as if it had a personal threshold, and the binary payoff is a modeling convenience rather than a fundamental restriction.
@@ D.1 Proof of the NP-hardness Result (Theorem˜2)
We reduce from the NP-complete Set Cover problem (Garey and Johnson, 2002): given a universe $U=\{u_{1},\dots,u_{n}\}$, a family of subsets $\mathcal{E}=\{E_{1},\dots,E_{r}\}$ with $\bigcup_{i\in[r]}E_{i}=U$, and an integer $m$, decide whether at most $m$ subsets cover $U$.
@@ D.2 Proof of the Approximation Guarantee (Theorem˜3)
In learning theory, for any distribution $\mathcal{D}$, a subset $\overline{\mathcal{H}}\subseteq\mathcal{H}$ is a $\gamma$-cover of $\mathcal{H}$ under $\mathcal{D}$ if, for every $h\in\mathcal{H}$, there exists $\overline{h}\in\overline{\mathcal{H}}$ such that $\mathcal{D}(\mathrm{DIS}(h,\overline{h}))\leq\gamma$. Haussler’s sphere-packing bound guarantees Haussler (1995) the existence of such a cover with